John is 2 times Bills age now. ten years ago he was seven times his age. What are their ages now.
Answer by stanbon(75887)

John is 2 times Bills age now. ten years ago he was seven times his age. What are their ages now.
------------------
Equations:
J = 2B
J-10 = 7(B-10)
---
Substitute for "J" and solve for "B":
2B-10 = 7B-70
5B = 60
B = 12 (Bill's age now)
-----
Since J = 2B, J = 2*12 = 24 (John's age now)
==============================================
Cheers,
Stan H.